OpenStudy (anonymous):

Solving quadratic equations by factoring Find the solution set of x^2 - 6x + 5 = 0 So far, I have (x^2 - 1x) (-5x + 5)

OpenStudy (anonymous):

I have a feeling I've set it up wrong already...

OpenStudy (jiteshmeghwal9):

First factor left side x^2-5x-x+5=0 x(x-5)-1(x-5)=0 Take (x-5) common so (x-1)(x-5)=0

OpenStudy (jiteshmeghwal9):

can u find x from here ?

OpenStudy (anonymous):

Yeah, my solution set would be {1,5} I was having trouble with the second set of parenthesis, I understand now. Thanks!

OpenStudy (jiteshmeghwal9):

yup & yw :)